# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ 5.2 Scattering of identical particles: quadrupole dominance
Next, we derive the energy loss rate for dSIDM processes involving identical particle scattering, which results in quadrupole emission. We use to evaluate (5.3), the energy differential cross-sections derived in Sec. 3.5.2, which take into account the $t$-channel, the $u$-channel, and their interference terms.
@@ 5.2.1 Massless emission
If the particle emitted is massless, the emission rate can be computed analytically using the energy differential cross-sections from Sec. 3.5.2 for $\kappa=0$.
@@
We start by considering processes of identical particle scattering mediated by a long-range force. In the case of vector emission, the processes $\chi\chi\to\chi\chi V$ and $SS\to SSV$ exhibit distinct energy differential cross-sections, leading to differences in their respective energy loss rates. In contrast, the processes $\chi\chi\to\chi\chi\phi$ and $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$ and $SS\to SS\phi$ and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$ yield, respectively, the same energy loss rate.
@@
$\dot{\epsilon}=\frac{\left(3\pi^{2}-20\right)n^{2}T^{3/2}}{m_{S}^{5/2}\pi^{7/2}}\times$ $\frac{5g^{6}}{72}$ for   $SS\to SSV$, (5.15a)
@@ 5.2.2 Massive emission
In the case of a massive emitted particle, the energy loss rate (5.1) has to be computed numerically using the energy differential cross-section from Sec. 3.5.2 for $\kappa\neq 0$.
@@
Figure 9: Effect of the mass of the emitted particle on the energy loss rate, for identical particle scattering, compared to a simple exponential suppression estimate [48] (black). Left panel: ratio $R$ (5.5) of the massive energy loss rate compared to the massless case, in the case of long-range mediation. Right panel: Comparison of the ratios $R_{j}$ in the different $j$-dSIDM scenarios to the benchmark result from the process $\chi\chi\to\chi\chi V$, under both long-range and short-range mediation. The limits $\left(A/m_{S}\right)^{2}\gg\lambda$ and $\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle\gg\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle$ (see App. C) are applied to the processes $SS\to SS\phi$ and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$. The orange region corresponds to $m_{\phi,V}\geq\mu\langle v_{i}^{2}\rangle$.
@@
Table 3: Ratios of the energy loss rates between the different dissipative scenarios, involving identical particle scattering mediated by a short-range interaction. Comparable DM candidate masses are assumed. The indices $k,l$ label the scenarios, with the shorthand notation listing the emitted particle first, followed by the DM candidates. The parameter $a^{4}a^{\prime 2}$ denotes the following configurations of coupling constants: $g^{2}g^{\prime 4}$ for $\chi\chi\to\chi\chi V$, $y^{2}y^{\prime 4}$ for $\chi\chi\to\chi\chi\phi$, $(A/m_{S})^{2}((A^{\prime}/m_{S})^{2}-{m_{\phi^{\prime}}^{2}\lambda_{S}}/2{m_{S}^{2}})^{2}$ for $SS\to SS\phi$, and $g^{2}(g^{\prime 2}+{m_{V^{\prime}}^{2}\lambda_{S}}/{8m_{S}^{2}})^{2}$ for $SS\to SSV$. First row: external-leg-dominated emission for $SS\to SS\phi$. Second row: vertex-dominated emission for $SS\to SS\phi$.
@@ 5.3 Soft emission
To conclude the discussion of the energy loss rate, we estimate, for long-range and short-range force mediated scattering, the contribution to the energy loss rates from soft emission using the soft energy-differential cross-sections derived in Sec. 4. We use the fiducial cut-off $\Lambda=\mu v_{i}^{2}/4$, such that particles emitted with an energy $\omega\lesssim\Lambda$ are considered soft.
@@ 5.3.1 Dipole soft emission
For dipole emission, the energy differential cross-section (4.12b) describes the $x\ll 1$ emission spectrum of the various dSIDM scenarios. For long-range mediation, substituting this expression into (5.3) and integrating over $\omega$ up to the cut-off $\Lambda$ yields the soft energy loss rate
$\left.\dot{\epsilon}_{\mathcal{D}}\right|_{\text{soft}}=f_{\phi,V}\mathcal{F}_{\mathcal{D}}^{2}\frac{n_{1}n_{2}\sqrt{\mu T}\left(1+\log 8\right)}{24\sqrt{2}\pi^{7/2}},$ (5.19)
where we added the subscript $\mathcal{D}$ to indicate the dipole nature of the emission. Of course, up to numerical prefactors, the rate exhibits the same parametric dependencies (temperature, mass, coupling constants) as the full result. Taking the ratio of this soft energy loss rate with the one obtained by integrating over the full energy range of dipole emission, i.e., the leading-order term in (5.6), denoted here as $\dot{\epsilon}_{\mathcal{D}}$, yields
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}\approx 0.76\leavevmode\nobreak\ \dot{\epsilon}_{\mathcal{D}}.$ (5.20)
@@
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}=\mathcal{F}_{\mathcal{D}}^{\prime 2}f_{\phi,V}\frac{2n_{1}n_{2}\sqrt{2}(\mu T)^{5/2}}{\pi^{7/2}m_{\phi^{\prime},V^{\prime}}^{4}}.$ (5.21)
@@
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}\approx 0.93\leavevmode\nobreak\ \dot{\epsilon}_{\mathcal{D}}.$ (5.22)
@@ 5.3.2 Quadrupole soft emission
We now turn to quadrupole emission, starting with long-range mediation. For distinguishable particle scattering, evaluating (5.3) up to the cutoff $\Lambda$ with the soft quadrupole differential cross-section (4.26), we find
@@
$\dot{\epsilon}_{\mathcal{Q}}|_{\text{soft}}=\mathcal{F}_{\mathcal{Q}}^{2}f_{\phi,V}\frac{n_{1}n_{2}(\mu T)^{3/2}\left[\alpha(1+\log 64)+\beta\right]}{15\sqrt{2}\pi^{7/2}}.$ (5.23)
@@
We added the subscript $\mathcal{Q}$ to indicate the quadrupole nature of the emission. The factor $\mathcal{F}_{\mathcal{Q}}$, along with the coefficients $\alpha$ and $\beta$, are defined in (3.29). We compare this soft energy loss rate to the full result obtained by integrating over the entire energy spectrum of quadrupole emission—i.e., the second term of (5.6), which we denote $\dot{\epsilon}_{\mathcal{Q}}$, in the first row of Tab. 4. For all the scenarios considered, between $80\%\text{ to }85\%$ of the total energy loss rate is carried by bosons with energies $\omega\lesssim\mu v_{i}^{2}/4$.
@@
For soft dissipative processes involving identical particles, the corresponding soft energy-loss rates are trivially obtained by inserting the energy differential cross-section (4.27a)–(4.29) into (5.3) and integrating up to the cut-off $\Lambda$. In the last row of Tab. 4, we compare these soft energy loss rates with those presented in (5.14a)-(5.15c). Relative to distinguishable-particle scattering, the soft-emission fraction is slightly larger for processes involving fermionic DM candidates and smaller for scalar DM candidates, the overall variation amounting to only a few percent. This is coming from the fact that in the $x\ll 1$ limit, some of the interference terms between the $t$- and $u$-channel are suppressed.
@@
We now proceed to the case of short-range mediating force. We evaluate (5.3) up to the cut-off $\Lambda$ using the soft quadrupole energy differential cross-section (4.25), and obtain
@@
$\dot{\epsilon}_{\mathcal{Q}}|_{\text{soft}}=\mathcal{F}_{\mathcal{Q}}^{\prime 2}f_{\phi,V}\frac{64\sqrt{2}n_{1}n_{2}(\mu T)^{7/2}\alpha}{15\pi^{7/2}m_{\phi^{\prime},V^{\prime}}^{4}},$ (5.24)
@@
where $\alpha$ and $\mathcal{F}_{\mathcal{Q}}^{\prime 2}$ are defined in (3.33)-(3.35). The ratios [^12] These ratios remain unchanged for identical particle scattering as both the soft and the full energy loss rates (5.16a)–(5.17b) acquire the same additional prefactors, which cancels in their quotient. of this soft expression compared to the full expression of the energy loss rate (second term in (5.8)) are derived in the second row of Tab. 4. As can be inferred from the behaviour of the soft energy differential cross-section in Fig. 5, the soft expression (5.24) being constant, overestimates the contribution of gauge bosons carrying energy close to $\Lambda$, yielding a ratio larger than unity. For $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$ and $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, the same approximation suppresses the non-trivial energy dependence of the full energy differential cross-section that becomes important near $\Lambda$, leading to a lower ratio compared to the other scenarios.
@@
Table 4: Comparison of the soft energy loss rate obtained by integrating the expression of the soft energy differential cross-section from Sec. 4 up to $\Lambda=\mu v_{i}^{2}/4$ with the full expressions presented in Sec. 5.1 (distinguishable particles) and Sec. 5.2 (identical particles), integrated to their kinematic endpoints.
@@ 6 Conclusions
Building on these amplitude-level results, we summarize the energy differential cross-sections in Tab. 5. Equipped with the latter, we compute in a final section the energy loss rates assuming that the DM candidates form an NR Maxwell–Boltzmann gas. An overview of the corresponding expressions is provided in Tab. 6.
@@ Acknowledgments.
Table 5: Summary of the Born energy-differential cross sections derived in this work. “Full spectrum”: expressions that include both hard and soft emissions. We present the massive-emission case (the massless limit follows by setting $\kappa=0$ or $C_{\phi,V}=0$). ”Soft spectrum”: results obtained by multiplying the elastic cross section by a soft factor (valid for $x\ll 1$). ”Additional radiation processes”: contributions from the dissipative channels Figs. 1(b)1(d).
@@
Table 6: Summary of the energy loss rates derived in this work, valid in the Born (perturbative) regime. “LO” denotes the first non-vanishing term in the temperature expansion of the energy loss rate, while “NLO” denotes the next-to-leading contribution. The rows labelled “Additional radiation processes” collect the contributions from the dissipative channel shown in Figs. 1(b)1(d). For these processes, results for distinguishable and identical particle scattering are reported on the same line.
@@ A.1 Center of Mass frame
$\omega=E_{1}+E_{2}-E_{3}-E_{4}\overset{\rm CM}{\simeq}\frac{\mu}{2}\left(|\vec{v}_{i}|^{2}-|\vec{v}_{f}|^{2}\right)+\mathcal{O}(v_{i}^{3})$ (A.2)
@@ A.2 Angular average
$\langle A\rangle_{\bm{\hat{}}{\mathbf{q}}}=\frac{1}{4\pi}\int d\Omega_{q}\,A,$ (A.4)
where $d\Omega_{q}$ is the solid-angle element centered around $\bm{\hat{}}{\mathbf{q}}$. Denoting by $\bm{\hat{}}{\mathbf{n}}_{j}$ any unit vector, we record the following useful identities used in this work:
@@ A.3 Feynman amplitudes
$i\mathcal{M}_{\chi_{1},\text{final}}$ $=\overline{u}(p_{3})\Gamma_{1}^{\mu}F(p_{3}+q)\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4})\overline{u}(p_{4})\Gamma^{\rho}_{2}u(p_{2})\epsilon^{*}_{\mu}(q),$ (A.6)
$i\mathcal{M}_{\chi_{2},\text{final}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4}-q)\overline{u}(p_{4})\Gamma^{\mu}_{2}F(p_{4}+q)\Gamma^{\rho}_{2}u(p_{2})\epsilon^{*}_{\mu}(q),$
$i\mathcal{M}_{\chi_{1},\text{initial}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}F\left(p_{1}-q\right)\Gamma^{\mu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4})\overline{u}(p_{4})\Gamma^{\rho}_{2}u(p_{2})\epsilon^{*}_{\mu}(q),$
$i\mathcal{M}_{\chi_{2},\text{initial}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4}-q)\overline{u}(p_{4})\Gamma^{\rho}_{2}F(p_{2}-q)\Gamma^{\mu}_{2}u(p_{2})\epsilon^{*}_{\mu}(q),$
@@
$i\mathcal{M}_{\tilde{S}_{1}}$ $=\Gamma^{\prime\mu\rho}_{1}\Delta_{\rho\nu}\left(p_{2}-p_{4}\right)\Gamma^{\nu}_{2}\epsilon^{*}_{\mu}(q),$ (A.7)
$i\mathcal{M}_{\tilde{S}_{2}}$ $=\Gamma_{1}^{\nu}\Delta_{\nu\rho}\left(p_{2}-p_{4}-q\right)\Gamma^{\prime\rho\mu}_{2}\epsilon^{*}_{\mu}(q),$

### v2

@@ 5.2 Scattering of identical particles: quadrupole dominance
Next, we derive the energy loss rate for dSIDM processes involving identical particle scattering, which results in quadrupole emission. We use to evaluate (5.3) the energy-differential cross sections derived in Sec. 3.5.2, which take into account the $t$-channel, the $u$-channel, and their interference terms.
@@ 5.2.1 Massless emission
If the particle emitted is massless, the emission rate can be computed analytically using the energy-differential cross sections from Sec. 3.5.2 for $\kappa=0$.
@@
We start by considering processes of identical particle scattering mediated by a long-range force. In the case of vector emission, the processes $\chi\chi\to\chi\chi V$ and $SS\to SSV$ exhibit distinct energy-differential cross sections, leading to differences in their respective energy loss rates. In contrast, the processes $\chi\chi\to\chi\chi\phi$ and $\tilde{\chi}\tilde{\chi}\to\tilde{\chi}\tilde{\chi}\phi$ and $SS\to SS\phi$ and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$ yield, respectively, the same energy loss rate.
@@
$\dot{\epsilon}=\frac{\left(3\pi^{2}-20\right)n^{2}T^{3/2}}{m_{S}^{5/2}\pi^{7/2}}\times$ $g^{6}\frac{5}{72}$ for   $SS\to SSV$, (5.15a)
@@ 5.2.2 Massive emission
In the case of a massive emitted particle, the energy loss rate (5.3) has to be computed numerically using the energy-differential cross section from Sec. 3.5.2 for $\kappa\neq 0$.
@@
Figure 9: Effect of the mass of the emitted particle on the energy loss rate, for identical particle scattering, compared to a simple exponential suppression estimate [54] (black). Left panel: ratio $R$ (5.5) of the massive energy loss rate compared to the massless case, for long-range mediation. Right panel: Comparison of the ratio $R_{j}$ in the different $j$-dSIDM scenarios to the benchmark result from the process $\chi\chi\to\chi\chi V$, under both long-range and short-range mediation. The limits $\left(A/m_{S}\right)^{2}\gg\lambda$ and $\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle\gg\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle$ (see App. C) are applied to the processes $SS\to SS\phi$ and $\tilde{S}\tilde{S}\to\tilde{S}\tilde{S}\phi$. The orange region corresponds to $m_{\phi,V}\geq\mu\langle v_{i}^{2}\rangle$.
@@
Table 3: Ratios of the energy loss rates between the different dissipative scenarios, involving identical particle scattering mediated by a short-range interaction. Comparable DM candidate masses are assumed. The indices $k,l$ label the scenarios, with the shorthand notation listing the emitted particle first, followed by the DM candidates. The parameter $a^{2}a^{\prime 4}$ denotes the following configurations of coupling constants: $g^{2}g^{\prime 4}$ for $\chi\chi\to\chi\chi V$, $y^{2}y^{\prime 4}$ for $\chi\chi\to\chi\chi\phi$, $(A/m_{S})^{2}((A^{\prime}/m_{S})^{2}-{m_{\phi^{\prime}}^{2}\lambda_{S}}/2{m_{S}^{2}})^{2}$ for $SS\to SS\phi$, and $g^{2}(g^{\prime 2}+{m_{V^{\prime}}^{2}\lambda_{S}}/{8m_{S}^{2}})^{2}$ for $SS\to SSV$. First row: external-leg-dominated emission for $SS\to SS\phi$. Second row: vertex-dominated emission for $SS\to SS\phi$.
@@ 5.3 Soft emission
To conclude the discussion of the energy loss rate, we estimate, for long-range and short-range force mediated scattering, the contribution to the energy loss rates from soft emission using the soft energy-differential cross sections derived in Sec. 4. We use the fiducial cut-off $\Lambda=\mu v_{i}^{2}/4$, such that particles emitted with an energy $\omega\lesssim\Lambda$ are considered soft.
@@ 5.3.1 Dipole soft emission
For dipole emission, the energy-differential cross sections (4.12b) describe the $x\ll 1$ emission spectrum of the various dSIDM scenarios. For long-range mediation, substituting the corresponding expression into (5.3) and integrating over $\omega$ up to the cut-off $\Lambda$ yields the soft energy loss rate
$\left.\dot{\epsilon}_{\mathcal{D}}\right|_{\text{soft}}=n_{1}n_{2}f_{\phi,V}\mathcal{F}_{\mathcal{D}}^{2}\frac{\left(1+\log 8\right)\sqrt{\mu T}}{24\sqrt{2}\pi^{7/2}},$ (5.19)
where we added the subscript $\mathcal{D}$ to label the dipole nature of the emission. Of course, up to numerical prefactors, the rate exhibits the same parametric dependencies (temperature, mass, coupling constants) as the full result. Taking the ratio of this soft energy loss rate with the one obtained by integrating over the full energy range of dipole emission, i.e., the leading-order term in (5.6), denoted here as $\dot{\epsilon}_{\mathcal{D}}$, yields
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}\approx 0.76\penalty 10000\ \dot{\epsilon}_{\mathcal{D}}.$ (5.20)
@@
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}=n_{1}n_{2}f_{\phi,V}\mathcal{F}_{\mathcal{D}}^{\prime 2}\frac{2\sqrt{2}(\mu T)^{5/2}}{\pi^{7/2}m_{\phi^{\prime},V^{\prime}}^{4}}.$ (5.21)
@@
$\dot{\epsilon}_{\mathcal{D}}|_{\text{soft}}\approx 0.93\penalty 10000\ \dot{\epsilon}_{\mathcal{D}}.$ (5.22)
@@ 5.3.2 Quadrupole soft emission
We now turn to quadrupole emission, starting with long-range mediation. For distinguishable particle scattering, evaluating (5.3) up to the cutoff $\Lambda$ with the soft quadrupole-differential cross section (4.26), we find
@@
$\dot{\epsilon}_{\mathcal{Q}}|_{\text{soft}}=n_{1}n_{2}f_{\phi,V}\mathcal{F}_{\mathcal{Q}}^{2}\frac{\left[\alpha(1+\log 64)+\beta\right](\mu T)^{3/2}}{15\sqrt{2}\pi^{7/2}}.$ (5.23)
@@
We added the subscript $\mathcal{Q}$ to label the quadrupole nature of the emission. The factor $\mathcal{F}_{\mathcal{Q}}$, along with the coefficients $\alpha$ and $\beta$, are defined in (3.34). We compare this soft energy loss rate to the full result obtained by integrating over the entire energy spectrum of quadrupole emission—i.e., the second term of (5.6), which we denote $\dot{\epsilon}_{\mathcal{Q}}$, in the first row of Tab. 4. For all the scenarios considered, between $80\%\text{ to }85\%$ of the total energy loss rate is carried by bosons with energies $\omega\lesssim\mu v_{i}^{2}/4$.
@@
For soft dissipative processes involving identical particles, the corresponding soft energy loss rates are trivially obtained by inserting the energy-differential cross sections (4.27)–(4.29) into (5.3) and integrating up to the cut-off $\Lambda$. In the last row of Tab. 4, we compare these soft energy loss rates with those presented in (5.14a)-(5.15b). Relative to distinguishable-particle scattering, the soft-emission fraction is slightly larger for processes involving fermionic DM candidates and smaller for scalar DM candidates, the overall variation amounting to only a few percent. This is coming from the fact that in the $x\ll 1$ limit, some of the interference terms between the $t$- and $u$-channel are suppressed.
@@
We now proceed to the case of short-range mediating force. We evaluate (5.3) up to the cut-off $\Lambda$ using the soft quadrupole energy-differential cross section (4.25), and obtain
@@
$\dot{\epsilon}_{\mathcal{Q}}|_{\text{soft}}=n_{1}n_{2}f_{\phi,V}\mathcal{F}_{\mathcal{Q}}^{\prime 2}\frac{64\sqrt{2}\alpha(\mu T)^{7/2}}{15\pi^{7/2}m_{\phi^{\prime},V^{\prime}}^{4}},$ (5.24)
@@
where $\alpha$ and $\mathcal{F}_{\mathcal{Q}}^{\prime 2}$ are defined in (3.38)-(3.40). The ratios [^13] These ratios remain unchanged for identical particle scattering as both the soft and the full energy loss rates (5.16a)–(5.17b) acquire the same additional prefactors, which cancels in their quotient. of this soft expression compared to the full expression of the energy loss rate (second term in (5.8)) are derived in the second row of Tab. 4. As can be inferred from the behaviour of the soft energy-differential cross section in Fig. 5, the soft expression (5.24) being constant, overestimates the contribution of gauge bosons carrying energy close to $\Lambda$, yielding a ratio larger than unity. For $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$ and $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, the same approximation suppresses the non-trivial energy dependence of the full energy-differential cross section that becomes important near $\Lambda$, leading to a lower ratio compared to the other scenarios.
@@
Table 4: Comparison of the soft energy loss rates obtained by integrating the expression of the soft energy-differential cross sections from Sec. 4 up to $\Lambda=\mu v_{i}^{2}/4$ with the full expressions presented in Sec. 5.1 (distinguishable particles) and Sec. 5.2 (identical particles), integrated to their kinematic endpoints.
@@ 6 Conclusions
Building on these amplitude-level results, we summarize the energy-differential cross sections in Tab. 5. Equipped with the latter, we compute in a final section the energy loss rates assuming that the DM candidates form an NR Maxwell–Boltzmann gas. An overview of the corresponding expressions is provided in Tab. 6.
@@ Acknowledgments.
Table 5: Summary of the Born energy-differential cross sections derived in this work. “Full spectrum”: expressions that include both hard and soft emissions. We present the massive-emission case (the massless limit follows by setting $\kappa=0$ or $C_{\phi,V}=0$). ”Soft spectrum”: results obtained by multiplying the elastic cross section by a soft factor (valid for $x\ll 1$). ”Additional radiation processes”: contributions from the dissipative channels Figs. 1(b)-1(d).
@@
Table 6: Summary of the energy loss rates derived in this work, valid in the Born (perturbative) regime. “LO” denotes the first non-vanishing term in the temperature expansion of the energy loss rate, while “NLO” denotes the next-to-leading contribution. The rows labelled “Additional radiation processes” collect the contributions from the dissipative channels shown in Figs. 1(b)-1(d). For these processes, results for distinguishable and identical particle scattering are reported on the same line.
@@ A.1 Center of Mass frame
$\omega=E_{1}+E_{2}-E_{3}-E_{4}\overset{\rm CM}{\simeq}\frac{\mu}{2}\left(|\vec{v}_{i}|^{2}-|\vec{v}_{f}|^{2}\right)+\mathcal{O}(v_{i}^{4}),$ (A.2)
@@ A.2 Angular average
$\langle A\rangle_{\bm{\hat{}}{\mathbf{q}}}=\frac{1}{4\pi}\int d\Omega_{\bm{\hat{}}{\mathbf{q}}}\,A,$ (A.4)
where $d\Omega_{\bm{\hat{}}{\mathbf{q}}}$ is the solid-angle element centered around $\bm{\hat{}}{\mathbf{q}}$. Denoting by $\bm{\hat{}}{\mathbf{n}}_{j}$ any unit vector, we record the following useful identities used in this work:
@@ A.3 Feynman amplitudes
$i\mathcal{M}_{\chi_{1},\text{final}}$ $=\overline{u}(p_{3})\Gamma_{1}^{\mu}F(p_{3}+q)\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4})\overline{u}(p_{4})\Gamma^{\rho}_{2}u(p_{2})\varepsilon^{*}_{\mu}(q),$ (A.6)
$i\mathcal{M}_{\chi_{2},\text{final}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4}-q)\overline{u}(p_{4})\Gamma^{\mu}_{2}F(p_{4}+q)\Gamma^{\rho}_{2}u(p_{2})\varepsilon^{*}_{\mu}(q),$
$i\mathcal{M}_{\chi_{1},\text{initial}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}F\left(p_{1}-q\right)\Gamma^{\mu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4})\overline{u}(p_{4})\Gamma^{\rho}_{2}u(p_{2})\varepsilon^{*}_{\mu}(q),$
$i\mathcal{M}_{\chi_{2},\text{initial}}$ $=\overline{u}(p_{3})\Gamma^{\nu}_{1}u(p_{1})\Delta_{\nu\rho}(p_{2}-p_{4}-q)\overline{u}(p_{4})\Gamma^{\rho}_{2}F(p_{2}-q)\Gamma^{\mu}_{2}u(p_{2})\varepsilon^{*}_{\mu}(q),$
@@
$i\mathcal{M}_{\tilde{S}_{1}}$ $=\Gamma^{\prime\mu\rho}_{1}\Delta_{\rho\nu}\left(p_{2}-p_{4}\right)\Gamma^{\nu}_{2}\varepsilon^{*}_{\mu}(q),$ (A.7)
$i\mathcal{M}_{\tilde{S}_{2}}$ $=\Gamma_{1}^{\nu}\Delta_{\nu\rho}\left(p_{2}-p_{4}-q\right)\Gamma^{\prime\rho\mu}_{2}\varepsilon^{*}_{\mu}(q),$